# Open-Source Multimodal Moxin Models with Moxin-VLM and Moxin-VLA

### v1

@@ 6.1 VLM Evaluation
We adopt the evaluation suit in Prismatic VLMs [karamcheti2024prismatic] to evaluate the VLM performance. It focuses on evaluations with well-defined metrics, spanning the following three areas:
Open-Ended Visual Question Answering. We evaluate on VizWiz [bigham2010vizwiz] and GQA [hudson2019gqa]. VizWiz assess general visual reasoning; VizWiz also contains a series of unanswerable questions. GQA evaluates spatial reasoning.
Localization. Part of the pretraining data mixture contains examples of predicting normalized bounding box coordinates given referring expressions in language. As such, we evaluate bounding box prediction accuracy on RefCOCO, RefCOCO+, and RefCOCOg [kazemzadeh2014referitgame], and on OCID-Ref [wang2021ocid]. RefCOCO focuses on short descriptions with spatial anchors, RefCOCO+ on strictly appearance based descriptions, and RefCOCOg on long, rich descriptions; OCID-Ref is a robotics dataset probing out-of-distribution generalization, with a focus on localizing objects in clutter.
Challenge Sets (Closed-Set Prediction). We evaluate on Visual Spatial Reasoning [liu2023visual], TallyQA [acharya2019tallyqa], and POPE [liu2023query]. VSR consists of challenging True/False questions about individual spatial relationships in diverse scenes; this is an especially challenging task, with most existing models failing to outperform the majority class baseline. TallyQA consists of questions that assess a VLM’s ability to count objects described in language, with expressions that range in complexity. POPE consists of targeted Yes/No questions that assess a VLM’s propensity to hallucinate.
@@ 6.2 VLA Evaluation
We evaluate the performance of Moxin-VLA within the LIBERO simulation environment [liu2023libero], aligning our experimental setup with the OpenVLA-OFT [kim2025fine] baseline to ensure a fair comparison. All models were fine-tuned on a cluster of 8 NVIDIA H100 GPUs using a batch size of 64 and a learning rate of 3e43e-4 . The model input consists of a history of two image frames combined with the robot’s proprioceptive state. To enable efficient adaptation, we utilized Low-Rank Adaptation (LoRA) with a rank of r=32.
@@
OpenVLA [kim2024openvla] 7.5B 84.7 88.4 79.2 53.7 76.5
SpatialVLA [qu2025spatialvla] 4.2B 88.2 89.9 78.6 55.5 78.1
CoT-VLA [zhao2025cot] 8.0B 87.5 91.6 87.6 69.0 81.1
NORA-Long [hung2025nora] 3.8B 92.2 95.4 89.4 74.6 87.9
@@
  • We evaluate the released OpenVLA-OFT [kim2025fine] checkpoints and observe large differences with the reported results in the paper, which is also mentioned in its Github issues. Thus we do not report OpenVLA-OFT here.

### v2

@@ 6.1 VLM Evaluation
We adopt the evaluation suit in Prismatic VLMs Karamcheti et al. (2024) to evaluate the VLM performance. It focuses on evaluations with well-defined metrics, spanning the following three areas:
Open-Ended Visual Question Answering. We evaluate on VizWiz Bigham et al. (2010) and GQA Hudson and Manning (2019). VizWiz assess general visual reasoning; VizWiz also contains a series of unanswerable questions. GQA evaluates spatial reasoning.
Localization. Part of the pretraining data mixture contains examples of predicting normalized bounding box coordinates given referring expressions in language. As such, we evaluate bounding box prediction accuracy on RefCOCO, RefCOCO+, and RefCOCOg Kazemzadeh et al. (2014), and on OCID-Ref Wang et al. (2021). RefCOCO focuses on short descriptions with spatial anchors, RefCOCO+ on strictly appearance based descriptions, and RefCOCOg on long, rich descriptions; OCID-Ref is a robotics dataset probing out-of-distribution generalization, with a focus on localizing objects in clutter.
Challenge Sets (Closed-Set Prediction). We evaluate on Visual Spatial Reasoning Liu et al. (2023b), TallyQA Acharya et al. (2019), and POPE Liu et al. (2023d). VSR consists of challenging True/False questions about individual spatial relationships in diverse scenes; this is an especially challenging task, with most existing models failing to outperform the majority class baseline. TallyQA consists of questions that assess a VLM’s ability to count objects described in language, with expressions that range in complexity. POPE consists of targeted Yes/No questions that assess a VLM’s propensity to hallucinate.
@@ 6.2 VLA Evaluation
We evaluate the performance of Moxin-VLA within the LIBERO simulation environment Liu et al. (2023a), aligning our experimental setup with the OpenVLA-OFT Kim et al. (2025) baseline to ensure a fair comparison. All models were fine-tuned on a cluster of 8 NVIDIA H100 GPUs using a batch size of 64 and a learning rate of 3e43e-4 . The model input consists of a history of two image frames combined with the robot’s proprioceptive state. To enable efficient adaptation, we utilized Low-Rank Adaptation (LoRA) with a rank of r=32.
@@
OpenVLA Kim et al. (2024) 7.5B 84.7 88.4 79.2 53.7 76.5
SpatialVLA Qu et al. (2025) 4.2B 88.2 89.9 78.6 55.5 78.1
CoT-VLA Zhao et al. (2025b) 8.0B 87.5 91.6 87.6 69.0 81.1
NORA-Long Hung et al. (2025) 3.8B 92.2 95.4 89.4 74.6 87.9
@@
  • We evaluate the released OpenVLA-OFT Kim et al. (2025) checkpoints and observe large differences with the reported results in the paper, which is also mentioned in its Github issues. Thus we do not report OpenVLA-OFT here.